# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

### v2

@@ III-A Feature Initialization
Recent LLIE methods demonstrate that preliminary enhancement facilitates subsequent restoration. Following Retinex theory [24], an image $I\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $L\in\mathbb{R}^{H\times W}$ as $I=R\odot L$, where $\odot$ denotes element-wise multiplication. Following EvLight [30], we enhance the low-light image $I$ through:
@@
where $\hat{L}$ is the estimated illumination map, implementing rough enhancement while preserving original details. The event stream $\{e_{k}\}$ is converted into a voxel grid $E\in\mathbb{R}^{H\times W\times B}$ ($B=32$) to preserve spatio-temporal information. The image $I_{lu}$ and event voxel $E$ are both downsampled to $\frac{H}{2}\times\frac{W}{2}$ via a two-layer $\times3$ conv with strides 1 & 2, extracting features $X^{\text{img}},X^{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ .
@@ III-B Cross-modal Feature Restoration
Existing methods for event-image fusion in low-light scenarios often fail to balance computational efficiency and robust cross-modal interaction, particularly in capturing high-frequency motion cues from events and preserving spatial details from images [17, 30]. Furthermore, they frequently treat event and image data as independent streams, failing to fully exploit the fine-grained spatiotemporal relationships between them. To address this, we propose the Cross-RWKV Module, which integrates Re-WKV Attention for capturing global interactions and Cross-Modal Switch Omni-Shift (CS-Shift) for local feature enhancement and multi-modal information interaction. Specifically, multiple Cross-RWKV Modules form a 4-level encoder-decoder. The encoder progressively downsamples spatial dimensions and expands channels, while the decoder upsamples to restore resolution, enabling multi-scale cross-modal fusion. The outputs $O_{\text{img}},O_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2C}$ is split by a 3×3 convolution into $\hat{X}_{\text{img}},\hat{X}_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$. As shown in Fig.4, the module operates through spatial and channel mix, enabling efficient cross-modal fusion of event and image features. The details of the Cross-RWKV Block are as follows:
@@ III-B1 Spatial Mix
where CS-shift is formulated as:
[…]
where $w_{i}^{img}$ and $w_{i}^{ev}$ represent learnable parameters for scaling specific branches, and $\text{Conv}_{k\times k}()$ denotes depthwise convolution with kernel size $k$. This CS-shift mechanism captures both local and global features by fusing information under different receptive fields, resulting in accurate token shift outcomes.
@@
Following the Restore-RWKV [57], we employ the Re-WKV mechanism to compute the global attention result $\text{wkv}\in\mathbb{R}^{T\times C}$. At the same time, key representations are directly exchanged between image tokens and event tokens, allowing each modality to attend to the spatial structure of the other modality, thereby enabling joint spatial reasoning between images and events:
@@
$\text{wkv}^{\text{ev}}=\text{Re-WKV}(K_{s}^{\text{img}},V_{s}^{\text{ev}})$ (8)
where Re-WKV is implemented by applying the Bi-WKV attention mechanism recurrently along both horizontal and vertical directions. Specifically, the attention output for the $t$-th token is computed as:
$\text{wkv}_{t}=\frac{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}+e^{u+k_{t}}}$ (9)
Here, $T$ denotes the total number of tokens, $k_{i}$ and $v_{i}$ are the key and value vectors at position $i$, and $w\in\mathbb{R}^{C}$, $u\in\mathbb{R}^{C}$ are learnable parameters that control the relative positional bias and provide additional weight to the current token, respectively. By applying Bi-WKV iteratively across multiple directions, Re-WKV effectively models long-range dependencies in 2D spatial structures, enabling enhanced global contextual reasoning across modalities.
@@
$X_{2}=\sigma(\alpha_{\text{evt}})\cdot\text{wkv}^{\text{ev}}+\left(1-\sigma(\alpha_{\text{evt}})\right)\cdot X^{\text{img}}$ (11)
@@
$O_{\text{evt}}=\left(\sigma(R_{\text{evt}})\odot X_{2}\right)W_{o}^{\text{evt}}$ (13)
W where $W_{o}^{\text{img}}$ and $W_{o}^{\text{evt}}$ are linear output projection matrices.
@@ III-B2 Channel Mix
$X_{\text{c}}^{\text{img}}=\text{OmniShift}(\text{Conv}(X_{\text{img}}))$ (14)
@@
$K_{\text{c}}=\text{squared ReLU}(X_{\text{c}}^{\text{img}}W_{k})$ (15)
@@
$V_{\text{c}}=K_{\text{c}}W_{v},\quad W_{v}\in\mathbb{R}^{C_{h}\times C}$ (16)
@@
$R_{\text{c}}=\sigma(X_{\text{c}}^{\text{img}}W_{r}),\quad W_{r}\in\mathbb{R}^{C\times C}$ (17)
@@
$O_{\text{c}}=R_{\text{c}}\odot V_{\text{c}}$ (18)
By restricting Channel Mix to the image modality only, we ensure more stable and informative feature transformations. Event streams, while rich in motion cues, often suffer from high sparsity and noise in the channel dimension, making them less suitable for dense feedforward modeling. Therefore, reserving event data for interaction in the spatial domain (via Spatial Mix) while using only image data for channel-wise transformations leads to a better division of modeling responsibilities, improving overall robustness and representation quality.
@@ III-C Frequency-aware Feature Fusion
TABLE I: Quantitative comparison of different methods on two real captured event datasets, SDE dataset and RELED dataset. The best and second best results are highlighted in bold and underlined respectively.
Method Publication Backbone Inputs SDE-in SDE-out RELED
[…]
SNR-Net[55] CVPR2022 Transformer $\checkmark$ $\times$ 20.05 0.630 22.18 0.661 26.47 0.851
Uformer[51] CVPR2022 Transformer $\checkmark$ $\times$ 21.09 0.752 22.32 0.725 28.86 0.832
LLFlow-L-SKF[53] CVPR2023 CNN $\checkmark$ $\times$ 20.92 0.661 21.68 0.647 28.62 0.862
Retinexformer[3] ICCV2023 Transformer $\checkmark$ $\times$ 21.3 0.692 22.92 0.683 26.66 0.865
RetinexMac[33] TCSVT2025 Transformer $\checkmark$ $\times$ 20.61 0.650 21.68 0.683 28.55 0.852
eSL-Net[46] ECCV2020 CNN $\checkmark$ $\checkmark$ 21.42 0.725 21.39 0.681 25.76 0.796
ELIE[17] TMM2023 Transformer $\checkmark$ $\checkmark$ 19.98 0.617 20.69 0.653 26.62 0.862
LLVE-SEG[35] AAAI2023 Transformer $\checkmark$ $\checkmark$ 21.79 0.705 22.35 0.690 29.19 0.875
ELEDNet[20] ECCV2024 Transformer $\checkmark$ $\checkmark$ 21.46 0.7126 22.91 0.726 31.96 0.910
EvLight[30] CVPR2024 Transformer $\checkmark$ $\checkmark$ 22.21 0.758 22.13 0.725 31.29 0.865
Ours - RWKV $\checkmark$ $\checkmark$ 23.09 0.766 23.43 0.767 32.18 0.910
@@ III-D Loss Function
We optimise EvRWKV in an end‑to‑end fashion by minimising a weighted combination of four complementary losses: the pixel‑wise reconstruction loss $\mathcal{L}_{r}$, the perceptual loss $\mathcal{L}_{p}$, the Structural Similarity loss $\mathcal{L}_{s}$, and its multi‑scale extension $\mathcal{L}_{m}$. Specifically, $\mathcal{L}_{\text{total}}$ is given by
@@
where the hyper‑parameters $\lambda_{r},\lambda_{p},\lambda_{s},\lambda_{m}$ balance the contributions of the individual terms.
To retain low‑level fidelity, we adopt the L1‑Charbonnier loss to penalise pixel errors between the enhanced image $I_{\mathrm{en}}$ and the ground‑truth image $I_{\mathrm{gt}}$:
$\mathcal{L}_{r}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(I_{\text{en}}-I_{\text{gt}})_{i}^{2}+\epsilon^{2}},$ (26)
where $I_{\mathrm{en}},I_{\mathrm{gt}}$ denote the enhanced and ground truth images and $\varepsilon=10^{-4}$ ensures numerical stability.
We employ the perceptual loss [25] to preserve the high-level similarity between the fused image and source images, which is given by
$\mathcal{L}_{p}=\bigl{\lVert}\,\Phi(I_{\mathrm{en}})-\Phi(I_{\mathrm{gt}})\,\bigr{\rVert}_{1},$ (27)
where $\Phi$ represents feature extraction using the Alex network [22].
For every $11\times 11$ Gaussian window $\Omega$ we compute the Structural Similarity (SSIM) index between the enhanced patch $x=I_{\mathrm{en}}\!\big{|}_{\Omega}$ and the corresponding ground-truth patch $y=I_{\mathrm{gt}}\!\big{|}_{\Omega}$:
$\operatorname{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (28)
where $\mu$, $\sigma$, and $\sigma_{xy}$ denote local means, standard deviations and cross-variance, while $C_{1}$ and $C_{2}$ avoid division by zero. Averaging the index over all windows and turning it into a loss gives
$\mathcal{L}_{s}=1-\operatorname{SSIM}(I_{\mathrm{en}},I_{\mathrm{gt}}).$ (29)
To capture structures at multiple resolutions, we build a five-level Gaussian pyramid and evaluate SSIM at every scale $j$. The resulting multi-scale similarity is the weighted geometric mean
$\operatorname{MS\text{-}SSIM}(x,y)=\prod_{j=1}^{5}\operatorname{SSIM}_{j}(x,y)^{\,\omega_{j}},$ (30)
where $\{\omega_{j}\}$ are the standard MS-SSIM weights ($\sum_{j}\omega_{j}=1$). The corresponding loss is obtained by
$\mathcal{L}_{m}=1-\operatorname{MS\text{-}SSIM}(I_{\mathrm{en}},I_{\mathrm{gt}}).$ (31)
@@ IV-A1 Datasets and Evaluation Metrics
SED [30], SDSD [47], and RELED [20] datasets are selected to test the enhancement performance of our EvRWKV framework.
SED dataset contains 91 image–event paired sequences (43 indoor and 48 outdoor) captured with a DAVIS346 event camera at a resolution of $346\times 260$; it is primarily used to evaluate the capability of different methods to leverage event cues for LLIE under diverse illumination conditions.
SDSD dataset provides 150 low-light/normal-light paired video sequences with an original resolution of $1920\times 1080$; following the dynamic-subset split (125 sequences for training and 25 for testing), all videos are down-sampled to $346\times 260$, and their event streams are synthesized with the v2e [15] simulator. This dataset evaluates the robustness of enhancement methods in dynamic low-light scenes.
@@
Figure 6: Visual comparisons with state-of-the-art methods on the SDE dataset.
Figure 7: Visual comparisons with state-of-the-art methods on the SDSD dataset.
Together, these three benchmarks provide a comprehensive testbed for assessing the multi-modal LLIE capability of our EvRWKV method. We use the peak-signal-to-noise ratio (PSNR) [14] and SSIM [52] for evaluation.
TABLE II: Quantitative comparison of different methods on the SDSD dataset. The best and second best results are highlighted in bold and underlined respectively.
Method Publication SDSD-in SDSD-out
$uparrow PSNR\uparrow SSIM\uparrow$
SNR-Net[55] CVPR’22 24.74 0.830 24.82 0.740
Uformer[51] CVPR’22 24.03 0.900 24.08 0.818
LLFlow-L-SKF[53] CVPR’23 23.39 0.818 20.39 0.634
Retinexformer[3] ICCV’23 25.90 0.852 26.08 0.815
RetinexMac[33] TCSVT’25 27.61 0.900 24.75 0.783
ELIE[17] TMM’23 27.46 0.879 23.29 0.742
eSL-Net[46] ECCV’20 23.68 0.821 24.95 0.805
LLVE-SEG[35] AAAI’23 27.58 0.888 23.51 0.726
ELEDNet[20] ECCV’24 28.52 0.912 25.08 0.828
EvLight[30] CVPR’24 28.49 0.909 25.17 0.817
Ours - 28.96 0.920 27.93 0.839

### v3

@@ III-A Feature Initialization
During data loading, we first apply gamma correction for initial global brightness adjustment. Recent LLIE methods demonstrate that preliminary enhancement facilitates subsequent restoration. Following Retinex theory [land1971lightness], an image $I\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $L\in\mathbb{R}^{H\times W}$ as $I=R\odot L$, where $\odot$ denotes element-wise multiplication. Beyond this global adjustment, following Retinexformer [cai2023retinexformer], we further enhance the gamma-corrected image $I$ through:
@@
where $\hat{L}$ is the estimated illumination map, implementing adaptive enhancement while preserving original details. The event stream $\{e_{k}\}$ is converted into a voxel grid $E\in\mathbb{R}^{H\times W\times B}$ ($B=32$) to preserve spatio-temporal information. The image $I_{lu}$ and event voxel $E$ are both downsampled to $\frac{H}{2}\times\frac{W}{2}$ via a two-layer $\times3$ convolution with strides 1 & 2, extracting features $X^{\text{img}},X^{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ .
@@ III-B Cross-modal Feature Restoration
Existing methods for event-image fusion in low-light scenarios often fail to balance computational efficiency and robust cross-modal interaction, particularly in capturing high-frequency motion cues from events and preserving spatial details from images [jiang2023event, liang2024towards]. Furthermore, they frequently treat event and image data as independent streams, failing to fully exploit the fine-grained spatiotemporal relationships between them. To address this, we propose the Cross-RWKV Module, which integrates Re-WKV [yang2025restore] Attention for capturing global interactions and Cross-Modal Switch Omni-Shift (CS-Shift) for local feature enhancement and multi-modal information interaction. Specifically, multiple Cross-RWKV Modules form a 4-level encoder-decoder. The encoder progressively downsamples spatial dimensions and expands channels, while the decoder upsamples to restore resolution, enabling multi-scale cross-modal fusion. The outputs $O_{\text{img}},O_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2C}$ is split by a 3×3 convolution into $\hat{X}_{\text{img}},\hat{X}_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$. As shown in Fig.4, the module operates through spatial and channel mix, enabling efficient cross-modal fusion of event and image features. The details of the Cross-RWKV Block are as follows:
@@ III-B1 Spatial Mix
where the CS-Shift is formulated as:
$X_{s}^{\text{img}}$ $=\sum_{i=1}^{4}w_{i}^{\text{img}}\cdot\text{Conv}_{k\times k}(X_{s}^{\text{img}}),$ (3)
$X_{s}^{\text{ev}}$ $=\sum_{i=1}^{4}w_{i}^{\text{ev}}\cdot\text{Conv}_{k\times k}(X_{s}^{\text{ev}}),$
where $w_{i}^{\text{img}}$ and $w_{i}^{\text{ev}}$ represent learnable parameters for scaling specific branches, and $\text{Conv}_{k\times k}()$ denotes depthwise convolution with kernel sizes of $1\times 1$, $3\times 3$, and $5\times 5$, respectively. The scaling weights are initialized with random normal distribution and updated via backpropagation during training. Each CS-Shift module is learned independently per stage and per modality. During test, the four parallel branches are reparameterized into a single $5\times 5$ convolution for efficiency. This CS-shift mechanism captures both local and global features by fusing information under different receptive fields, resulting in accurate token shift outcomes.
@@
Following the Restore-RWKV [yang2025restore], we employ the Re-WKV mechanism to compute the global attention result $\text{wkv}\in\mathbb{R}^{T\times C}$. At the same time, key representations are directly exchanged between image tokens and event tokens, allowing each modality to attend to the spatial structure of the other modality, thereby enabling joint spatial reasoning between images and events:
@@
$\text{wkv}^{\text{ev}}=\text{Re-WKV }(K_{s}^{\text{img}},V_{s}^{\text{ev}}),$ (8)
where Re-WKV [yang2025restore] is implemented by recurrently applying Bidirectional (Bi-WKV) attention along alternating scan directions. As shown in Fig. 5, the recurrent process can be formulated as:
$\text{wkv}^{(j)}=\text{Bi-WKV}^{(j)}(\Delta_{\text{dir}}(K_{s}),\Delta_{\text{dir}}(\text{wkv}^{(j-1)})),$ (9)
where $\text{Bi-WKV}^{(j)}(\cdot)$ denotes the $j$-th Bi-WKV attention, $\Delta_{\text{dir}}(\cdot)$ represents the direction-changing operation that alternates between horizontal scan (H-Scan) and vertical scan (V-Scan), and $\text{wkv}^{(0)}=V_{s}$. The final Re-WKV output is $\text{wkv}^{(2)}$. Specifically, the Bi-WKV attention output for the $t$-th token is computed as:
$\text{wkv}_{t}=\frac{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}+e^{u+k_{t}}}.$ (10)
Figure 5: Illustration of Re-WKV mechanism. Bi-WKV attention is applied recurrently along alternating horizontal and vertical scan directions, with each iteration taking the previous attention result as input.
Here, $T$ denotes the total number of tokens, $k_{i}$ and $v_{i}$ are the key and value vectors at position $i$, and $w\in\mathbb{R}^{C}$, $u\in\mathbb{R}^{C}$ are learnable parameters that control the relative positional bias and provide additional weight to the current token, respectively. By applying Bi-WKV iteratively across multiple directions, Re-WKV [yang2025restore] effectively models long-range dependencies in 2D spatial structures, enabling enhanced global contextual reasoning across modalities.
@@
$X_{2}=\sigma(\alpha_{\text{evt}})\cdot\text{wkv}^{\text{ev}}+\left(1-\sigma(\alpha_{\text{evt}})\right)\cdot X^{\text{img}}.$ (12)
@@
$O_{\text{evt}}=\left(\sigma(R_{\text{evt}})\odot X_{2}\right)W_{o}^{\text{evt}},$ (14)
where $W_{o}^{\text{img}}$ and $W_{o}^{\text{evt}}$ are linear output projection matrices.
@@ III-B2 Channel Mix
$X_{\text{c}}^{\text{img}}=\text{OmniShift}(\text{Conv}(X_{\text{img}})).$ (15)
@@
$K_{\text{c}}=\text{squared ReLU}(X_{\text{c}}^{\text{img}}W_{k}),$ (16)
@@
$V_{\text{c}}=K_{\text{c}}W_{v},\quad W_{v}\in\mathbb{R}^{C_{h}\times C}.$ (17)
@@
$R_{\text{c}}=\sigma(X_{\text{c}}^{\text{img}}W_{r}),\quad W_{r}\in\mathbb{R}^{C\times C}.$ (18)
@@
$O_{\text{c}}=R_{\text{c}}\odot V_{\text{c}}.$ (19)
By restricting Channel Mix to the image modality only, we ensure more stable and informative feature transformations. Event streams, while rich in motion cues, often suffer from high sparsity and noise in the channel dimension, making them less suitable for dense feedforward modeling. Therefore, reserving event data for interaction in the spatial domain while using only image data for channel-wise transformations leads to a better division of modeling responsibilities, improving overall robustness and representation quality.
@@ III-D Loss Function
The loss function $\mathcal{L}$ used for training our EvRWKV model is composed of four complementary terms, formulated as:
@@
where $\lambda_{r}=1$, $\lambda_{p}=0.8$, $\lambda_{s}=1$, and $\lambda_{m}=0.5$ are balancing weights. The reconstruction term combines pixel-wise fidelity and perceptual similarity:
$\mathcal{L}_{r}=\sqrt{(I_{\text{en}}-I_{\text{gt}})_{i}^{2}+\epsilon^{2}},$ (27)
$\mathcal{L}_{p}=\|\Phi(I_{\text{en}})-\Phi(I_{\text{gt}})\|_{1},$
where $\epsilon=10^{-4}$ and $\Phi$ denoting AlexNet-based feature extraction.
To ensure enhanced images maintain natural structural characteristics and visual quality, we employ both the Structural Similarity (SSIM) loss and its multi-scale extension (MS-SSIM):
$\mathcal{L}_{s}=1-\text{SSIM}(I_{\text{en}},I_{\text{gt}}),$ (28)
$\mathcal{L}_{m}=1-\text{MS-SSIM}(I_{\text{en}},I_{\text{gt}}).$
TABLE I: Quantitative comparison of different methods on the SDE dataset. The best and second-best results are highlighted in bold and underlined, respectively.
  Method Backbone Inputs SDE-in SDE-out
[…]
SNR-Net [xu2022snr] (CVPR’22) Transformer 20.05 0.630 0.244 14.86 22.18 0.661 0.184 12.16
Uformer [wang2022uformer] (CVPR’22) Transformer 21.09 0.752 0.109 9.670 22.32 0.747 0.097 8.969
LLFlow [wu2023learning] (CVPR’23) CNN 20.92 0.661 0.225 11.76 21.68 0.647 0.236 14.56
Retinexformer [cai2023retinexformer] (ICCV’23) Transformer 21.30 0.692 0.124 11.24 22.92 0.683 0.146 11.35
RetinexMac [10810469] (TCSVT’25) Transformer 20.61 0.650 0.149 12.54 21.68 0.683 0.170 11.12
eSL-Net [wang2020event] (ECCV’20) CNN 21.42 0.725 0.130 10.62 21.39 0.681 0.282 15.81
ELIE [jiang2023event] (TMM’23) Transformer 19.98 0.617 0.217 14.75 20.69 0.653 0.313 15.98
LLVE-SEG [liu2023low] (AAAI’23) Transformer 21.79 0.705 0.114 10.54 22.35 0.690 0.154 12.39
ELEDNet [kim2024towards] (ECCV’24) Transformer 21.46 0.713 0.116 13.60 22.91 0.726 0.129 12.77
EvLight [liang2024towards] (CVPR’24) Transformer 22.21 0.758 0.101 11.13 22.13 0.725 0.090 11.48
Ours RWKV 23.09 0.770 0.094 8.534 23.43 0.768 0.091 8.926
TABLE II: Quantitative comparison of different methods on the SDSD and RELED datasets. The best and second-best results are highlighted in bold and underlined, respectively.
  Method SDSD-in SDSD-out RELED
[…]
SNR-Net [xu2022snr] 24.74 0.830 0.097 7.963 24.82 0.740 0.107 4.638 26.47 0.851 0.192 14.76
Uformer [wang2022uformer] 26.60 0.881 0.068 6.607 24.08 0.818 0.137 4.773 28.86 0.832 0.212 14.40
LLFlow [wu2023learning] 23.39 0.818 0.104 9.586 20.39 0.634 0.255 7.468 28.62 0.862 0.154 13.07
Retinexformer [cai2023retinexformer] 25.90 0.852 0.086 6.954 26.08 0.815 0.061 4.322 26.66 0.865 0.168 13.54
RetinexMac [10810469] 27.61 0.900 0.046 6.550 24.75 0.783 0.103 5.372 28.55 0.851 0.192 12.26
eSL-Net [wang2020event] 23.68 0.821 0.067 7.299 24.95 0.805 0.063 4.821 25.76 0.796 0.312 16.11
ELIE [jiang2023event] 27.46 0.879 0.074 6.736 23.29 0.742 0.116 5.568 26.62 0.862 0.267 14.14
LLVE-SEG [liu2023low] 27.58 0.888 0.053 6.837 23.51 0.726 0.110 5.036 29.19 0.875 0.098 11.94
ELEDNet [kim2024towards] 28.50 0.910 0.040 7.439 25.17 0.817 0.081 4.895 31.96 0.910 0.106 10.82
EvLight [liang2024towards] 28.52 0.912 0.039 6.764 25.08 0.828 0.060 4.455 31.29 0.885 0.067 10.63
Ours 28.96 0.920 0.032 6.451 27.93 0.839 0.058 4.110 32.18 0.910 0.044 9.207
@@ IV-A1 Datasets and Evaluation Metrics
SDE [liang2024towards], SDSD [wang2021seeing], and RELED [kim2024towards] datasets are selected to test the enhancement performance of our EvRWKV framework.
SDE dataset contains 91 image–event paired sequences (43 indoor and 48 outdoor) captured with a DAVIS346 event camera at a resolution of $346\times 260$; it is primarily used to evaluate the capability of different methods to leverage event cues for LLIE under diverse illumination conditions.
SDSD dataset provides 150 low-light/normal-light paired video sequences with an original resolution of $1920\times 1080$; following the dynamic-subset split (125 sequences for training and 25 for testing), all videos are down-sampled to $346\times 260$, and their event streams are synthesized with the v2e [hu2021v2e] simulator. This dataset evaluates the robustness of enhancement methods in dynamic low-light scenes.
@@
Figure 7: Visual comparisons with state-of-the-art methods on SDE-in and SDE-out dataset [liang2024towards].
Figure 8: Visual comparisons with state-of-the-art methods on SDSD-in and SDSD-out dataset [wang2021seeing].
Together, these three benchmarks provide a comprehensive testbed for assessing the multi-modal LLIE capability of our EvRWKV method. We use the peak-signal-to-noise ratio (PSNR) [hore2010image], structural similarity index (SSIM) [wang2004image], learned perceptual image patch similarity (LPIPS) [zhang2018unreasonable], and natural image quality evaluator (NIQE) [mittal2012making] for quantitative evaluation, where higher PSNR and SSIM values and lower LPIPS and NIQE scores indicate better performance.
Figure 9: Visual comparisons with state-of-the-art methods on the RELED dataset [kim2024towards].
Figure 10: Cross-dataset generalization results on the LIE [jiang2023event] dataset using the model trained on SDE-out.
TABLE III: Cross-dataset generalization on LIE (trained on SDE-out). The best and second-best results are in bold and underlined.
  Method Uformer [wang2022uformer] RetinexMac [10810469] ELEDNet [kim2024towards] EvLight [liang2024towards] Ours
[…]